OpenStudy (anonymous):

Alex stocks up for winter. He buys 32 cans of vegetables. He pays 80 cents per can for tomatoes and 40 cents per can for corn, for a total cost of $18.00. How many cans of corn does he buy?

OpenStudy (anonymous):

80x+40y=1800 80x+40y=32 get x alone: 80x=32-40y x=(32-40y)/80 plug it in to solve for y: 80{(32-40y)/80}+40y=32 I don't know what to do after that o.o